# On the Fairness of Privacy Protection: Measuring and Mitigating the Disparity of Group Privacy Risks for Differentially Private Machine Learning

### v1

@@ 5.1 Experimental setup
Full experimental details are provided in App. D.1. We conduct experiments on datasets commonly used in privacy and fairness research [2, 3, 27], including MNIST [14], as well as three fairness-related datasets: two tabular datasets, Adult and Law[13], and one image dataset, UTKFace[33]. Our study compares three training algorithms: standard SGD, DP-SGD, and our proposed DP-SGD-S. For both DP-SGD and DP-SGD-S, the default privacy parameters are set to $(\epsilon,\delta)=(10,1\mathrm{e}{-5})$, and the default scale bound $\tau$ for DP-SGD-S is set to 2. Three model architectures are considered: LR, MLP, and CNN. To measure fairness in privacy protection across groups, we use the GPRP metric, assessed via our auditing method PA-ALOOA. The model utility is evaluated through classification accuracy. All results reported represent the average of five independent runs, with all values presented in percentage points for clarity.
@@ A.1 DP-SGD
In our work, we concentrate on DP-SGD to uphold model privacy. Building upon SGD, the fundamental method for training a model $f$ with parameters $w$ by minimizing the empirical loss function $\ell(\hat{y},y)$ for prediction $\hat{y}$ and label $y$, DP-SGD (as illustrated in Algo. 2) integrates gradient clipping and noise addition for achieving the $(\epsilon,\delta)$-DP guarantees. In Algo. 2, during each epoch, per-sample gradients $g_{i}$ are computed (Line 5). Since these gradients typically have unbounded sensitivity, they are clipped to ensure their norm does not exceed the hyperparameter $C$ (Line 6). The clipped gradients are then aggregated and Gaussian noise is added to yield $\tilde{g}$ (Line 8). $\tilde{g}$ is subsequently scaled by the learning rate $\eta$ and utilized for parameter update (Line 9). The final accumulated $(\epsilon,\delta)$, which is calculated by Rényi differential privacy (RDP) [16] and the moment accounting mechanism proposed by [1], quantifies the privacy protection ability.
Algorithm 2 DP-SGD [1]
@@ A.2 Definitions of Membership Inference Games
Definition 6 (Average-case Membership Inference Game [30]).
@@
Definition 7 (Worst-case Membership Inference Game [29]).
@@ A.3 Privacy Audting by Different Attacks
We introduce existing algorithms that use average-case attacks for privacy auditing (PA-ACAs). Specifically, one approach conducts privacy auditing using the global attack (PA-GA) [28], which is detailed in Algo. 3. In Algo. 3, a single threshold $\beta$ is determined based on the overall behavior of all auditing samples. Another approach performs privacy auditing through the group-based attack (PA-GBA) [4], as described in Algo. 4. PA-GBA provides the adversary with background knowledge about which group $g_{i}$ each data point $(\mathbf{x}_{i},y_{i})$ belongs to. It then determines $K$ thresholds $\beta^{k}$ based on the behavior of all auditing samples within each group, where $K$ represents the number of groups. As illustrated in Algo. 3 and Algo. 4, a single execution of the attack generates a prediction for the membership status of each data point in the auditing dataset. To evaluate the privacy risk associated with individual data points more comprehensively, researchers typically conduct multiple iterations of privacy auditing using average-case attacks [4]. Each iteration yields new results, which are aggregated to estimate the likelihood of a data point being accurately identified as either a member or a non-member.
Algorithm 3 PA-GA [28]
@@
In recent years, numerous studies have focused on using privacy auditing to evaluate the differential privacy (DP) guarantees of the DP-SGD algorithm [18, 23, 32, 12, 2, 17]. These studies aim to bridge the gap between theoretical guarantees and practical performance, offering empirical insights into the actual privacy leakage in real-world deployments. A common approach in these studies is Privacy Auditing via the Leave-One-Out Attack (PA-LOOA), as outlined in Algo. 5. The algorithm iteratively assesses the impact of including or excluding a specific data record zz—often crafted as a worst-case scenario for auditing DP-SGD—within the training dataset $D$ (Lines 2–8). For each repetition, the framework trains two models: $f_{0}$, using the modified dataset $D\setminus z$, and $f_{1}$, using the original dataset $D$ (Lines 3–4). The outputs of these models are recorded, and the membership status of the data record is tracked (Lines 5–7). Based on these outputs and the membership status, attack scores are computed to estimate the likelihood of the record’s inclusion (Line 9). Finally, assuming an optimal adversary conducting the attack, an optimal threshold is applied to infer whether the record $z$ was part of the training dataset (Line 10).
@@
Algorithm 5 PA-LOOA [18, 23, 32, 12, 2, 17]
@@ D.1 Details of Experimental Setup
The MNIST dataset [14] comprises 60,000 training and 10,000 testing samples, with each sample being a 28×28 grayscale image of a handwritten digit from 0 to 9, spanning ten classes. Due to computational limitations and to improve training efficiency, we randomly select 1,000 samples per class from the original dataset, resulting in a balanced dataset of 10,000 samples used in our experiments. In our study, we treat the classification label (i.e., the digit class) as a proxy for the demographic group to analyze group-specific privacy risks.
For the tabular fairness-related datasets, Adult, Bank, Credit, and Law [13] [^1] Dataset can be download from https://github.com/tailequy/fairness_dataset/tree/main/experiments/data, the detailed information is shown in Tab. 3. For the image-based fairness-related dataset, UTKFace [33], we conduct evaluation after data cleaning and preprocessing. [^2] Dataset can be downloaded from https://www.kaggle.com/datasets/nipunarora8/age-gender-and-ethnicity-face-data-csv/data?select=age_gender.csv The final dataset consists of 27,305 grayscale facial images of size 1×48×48. In our experiments, we treat ethnicity as the protected attribute and gender as the prediction label. The ethnicity attribute includes five classes, with the number of samples per class being 10,078; 4,526; 3,434; 3,975; and 1,692, respectively.
Table 3: The information of experimental datasets. Here, 0 represents the advantaged group, and 11 represents the disadvantaged group.
@@
Our study compares three training algorithms: standard SGD, DP-SGD, and our proposed DP-SGD-S. DPML algorithms (i.e., DP-SGD and DP-SGD-S) are implemented using the Opacus library [31]. For the SGD algorithm, all datasets are trained using the SGD optimizer with a learning rate of 0.1. For DPML algorithms, the privacy hyperparameters $(\epsilon,\delta)$ are configured with $(10,1\mathrm{e}{-5})$ and $(1,1\mathrm{e}{-5})$. For DP-SGD-S, the default scale bound $\tau$ is set to 2. When $\epsilon=10$, all datasets use the SGD optimizer with a learning rate of 0.1, and the clipping bound is set to 10. When $\epsilon=1$, the tabular datasets use the same optimizer and learning rate as above, while the image datasets are trained using the Adam optimizer with a learning rate of 0.005. In this case, the gradient clipping bound is set to 5 for all datasets. Across all experiments, we use a batch size of 256 and train for 20 epochs.
@@ D.2.4 The results of outcome fairness
Traditional definitions of outcome fairness are primarily designed for binary classification tasks where the sensitive attribute has only two groups (e.g., male vs. female) [19]. Under this setting, fairness metrics are typically computed using the absolute difference between the values of the two groups. However, in the case of the UTKFace dataset, the sensitive attribute is multi-class, containing more than two categories. To accommodate this, we adopt a natural extension of the conventional definitions: instead of calculating the absolute difference between two groups, we compute the range across all subgroups, i.e., the difference between the maximum and minimum values of the fairness metric across the group set. This allows us to quantify disparities in outcomes among multiple demographic groups.

### v3

@@ 5.1 Experimental setup
Full experimental details are provided in App. D.1. We conduct experiments on datasets commonly used in privacy and fairness research [annamalai2024nearly, bagdasaryan2019differential, xu2021removing], including MNIST [lecun1998mnist], as well as three fairness-related datasets: two tabular datasets, Adult and Law[le2022survey], and one image dataset, UTKFace[zhang2017age]. Our study compares three training algorithms: standard SGD, DP-SGD, and our proposed DP-SGD-S. For both DP-SGD and DP-SGD-S, the default privacy parameters are set to $(\epsilon,\delta)=(10,1\mathrm{e}{-5})$, and the default scale bound $\tau$ for DP-SGD-S is set to 2. Three model architectures are considered: LR, MLP, and CNN. To measure fairness in privacy protection across groups, we use the GPRP metric, assessed via our auditing method PA-ALOOA. The model utility is evaluated through classification accuracy. All results reported represent the average of five independent runs, with all values presented in percentage points for clarity.
@@ A.1 DP-SGD
In our work, we concentrate on DP-SGD to uphold model privacy. Building upon SGD, the fundamental method for training a model $f$ with parameters $w$ by minimizing the empirical loss function $\ell(\hat{y},y)$ for prediction $\hat{y}$ and label $y$, DP-SGD (as illustrated in Algo. 2) integrates gradient clipping and noise addition for achieving the $(\epsilon,\delta)$-DP guarantees. In Algo. 2, during each epoch, per-sample gradients $g_{i}$ are computed (Line 5). Since these gradients typically have unbounded sensitivity, they are clipped to ensure their norm does not exceed the hyperparameter $C$ (Line 6). The clipped gradients are then aggregated and Gaussian noise is added to yield $\tilde{g}$ (Line 8). $\tilde{g}$ is subsequently scaled by the learning rate $\eta$ and utilized for parameter update (Line 9). The final accumulated $(\epsilon,\delta)$, which is calculated by Rényi differential privacy (RDP) [mironov2017renyi] and the moment accounting mechanism proposed by [abadi2016deep], quantifies the privacy protection ability.
Algorithm 2 DP-SGD [abadi2016deep]
@@ A.2 Definitions of Membership Inference Games
Definition 6 (Average-case Membership Inference Game [yeom2018privacy]).
@@
Definition 7 (Worst-case Membership Inference Game [ye2022enhanced]).
@@ A.3 Privacy Audting by Different Attacks
We introduce existing algorithms that use average-case attacks for privacy auditing (PA-ACAs). Specifically, one approach conducts privacy auditing using the global attack (PA-GA) [yaghini2022disparate], which is detailed in Algo. 3. In Algo. 3, a single threshold $\beta$ is determined based on the overall behavior of all auditing samples. Another approach performs privacy auditing through the group-based attack (PA-GBA) [chang2021privacy], as described in Algo. 4. PA-GBA provides the adversary with background knowledge about which group $g_{i}$ each data point $(\mathbf{x}_{i},y_{i})$ belongs to. It then determines $K$ thresholds $\beta^{k}$ based on the behavior of all auditing samples within each group, where $K$ represents the number of groups. As illustrated in Algo. 3 and Algo. 4, a single execution of the attack generates a prediction for the membership status of each data point in the auditing dataset. To evaluate the privacy risk associated with individual data points more comprehensively, researchers typically conduct multiple iterations of privacy auditing using average-case attacks [chang2021privacy]. Each iteration yields new results, which are aggregated to estimate the likelihood of a data point being accurately identified as either a member or a non-member.
Algorithm 3 PA-GA [yaghini2022disparate]
@@
In recent years, numerous studies have focused on using privacy auditing to evaluate the differential privacy (DP) guarantees of the DP-SGD algorithm [nasr2023tight, steinke2023privacy, zanella2023bayesian, jagielski2020auditing, annamalai2024nearly, nasr2021adversary]. These studies aim to bridge the gap between theoretical guarantees and practical performance, offering empirical insights into the actual privacy leakage in real-world deployments. A common approach in these studies is Privacy Auditing via the Leave-One-Out Attack (PA-LOOA), as outlined in Algo. 5. The algorithm iteratively assesses the impact of including or excluding a specific data record zz—often crafted as a worst-case scenario for auditing DP-SGD—within the training dataset $D$ (Lines 2–8). For each repetition, the framework trains two models: $f_{0}$, using the modified dataset $D\setminus z$, and $f_{1}$, using the original dataset $D$ (Lines 3–4). The outputs of these models are recorded, and the membership status of the data record is tracked (Lines 5–7). Based on these outputs and the membership status, attack scores are computed to estimate the likelihood of the record’s inclusion (Line 9). Finally, assuming an optimal adversary conducting the attack, an optimal threshold is applied to infer whether the record $z$ was part of the training dataset (Line 10).
@@
Algorithm 5 PA-LOOA [nasr2023tight, steinke2023privacy, zanella2023bayesian, jagielski2020auditing, annamalai2024nearly, nasr2021adversary]
@@ D.1 Details of Experimental Setup
The MNIST dataset [lecun1998mnist] comprises 60,000 training and 10,000 testing samples, with each sample being a 28×28 grayscale image of a handwritten digit from 0 to 9, spanning ten classes. Due to computational limitations and to improve training efficiency, we randomly select 1,000 samples per class from the original dataset, resulting in a balanced dataset of 10,000 samples used in our experiments. In our study, we treat the classification label (i.e., the digit class) as a proxy for the demographic group to analyze group-specific privacy risks.
For the tabular fairness-related datasets, Adult, Bank, Credit, and Law [le2022survey] [^1] Dataset can be download from https://github.com/tailequy/fairness_dataset/tree/main/experiments/data, the detailed information is shown in Tab. 3. For the image-based fairness-related dataset, UTKFace [zhang2017age], we conduct evaluation after data cleaning and preprocessing. [^2] Dataset can be downloaded from https://www.kaggle.com/datasets/nipunarora8/age-gender-and-ethnicity-face-data-csv/data?select=age_gender.csv The final dataset consists of 27,305 grayscale facial images of size 1×48×48. In our experiments, we treat ethnicity as the protected attribute and gender as the prediction label. The ethnicity attribute includes five classes, with the number of samples per class being 10,078; 4,526; 3,434; 3,975; and 1,692, respectively.
Table 3: The information of experimental datasets. Here, 0 represents the advantaged group, and $1$ represents the disadvantaged group.
@@
Our study compares three training algorithms: standard SGD, DP-SGD, and our proposed DP-SGD-S. DPML algorithms (i.e., DP-SGD and DP-SGD-S) are implemented using the Opacus library [opacus]. For the SGD algorithm, all datasets are trained using the SGD optimizer with a learning rate of 0.1. For DPML algorithms, the privacy hyperparameters $(\epsilon,\delta)$ are configured with $(10,1\mathrm{e}{-5})$ and $(1,1\mathrm{e}{-5})$. For DP-SGD-S, the default scale bound $\tau$ is set to 2. When $\epsilon=10$, all datasets use the SGD optimizer with a learning rate of 0.1, and the clipping bound is set to 10. When $\epsilon=1$, the tabular datasets use the same optimizer and learning rate as above, while the image datasets are trained using the Adam optimizer with a learning rate of 0.005. In this case, the gradient clipping bound is set to 5 for all datasets. Across all experiments, we use a batch size of 256 and train for 20 epochs.
@@ D.2.4 The results of outcome fairness
Traditional definitions of outcome fairness are primarily designed for binary classification tasks where the sensitive attribute has only two groups (e.g., male vs. female) [pessach2022review]. Under this setting, fairness metrics are typically computed using the absolute difference between the values of the two groups. However, in the case of the UTKFace dataset, the sensitive attribute is multi-class, containing more than two categories. To accommodate this, we adopt a natural extension of the conventional definitions: instead of calculating the absolute difference between two groups, we compute the range across all subgroups, i.e., the difference between the maximum and minimum values of the fairness metric across the group set. This allows us to quantify disparities in outcomes among multiple demographic groups.